OpenStudy (anonymous):

graph the equation and identify the y-concept y=1/4x

OpenStudy (anonymous):

y intercept is at the origin. the graph: http://bit.ly/peazsg

OpenStudy (anonymous):

I dont know what that means

OpenStudy (anonymous):

you don't know what what means?

OpenStudy (anonymous):

origin is (0,0). y-intercept is the place where the graph crosses the y axis. The link is to the graph of y=1/4x

OpenStudy (anonymous):

actually, I don't know if you meant y=(1/4)x or y=1/(4x). I assumed the first one...

OpenStudy (anonymous):

so (0,0) is the y-intercept?

OpenStudy (anonymous):

yes.